OpenStudy (thickgirl_angel):

Find the 4th term if the sequend in which a1 = 2 and an+1 = -4an + 2.

jhonyy9 (jhonyy9):

so than you know this formula a_(n+1) = -4an +2 and you know that a_1 = 2so how you get the a_2 ?

jhonyy9 (jhonyy9):

ok. a little help you know that a_1 = 2 so than you can writing for a_2 : a_2 = -4a_1 +2 = -4*2 +2 = -8 +2 = -6 right ?

jhonyy9 (jhonyy9):

can you calculi now using these above wrote the a_3 = ?

jhonyy9 (jhonyy9):

a_3 = -4*a_2 +2 = ?

jhonyy9 (jhonyy9):

please collaborate

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (thickgirl_angel):

Yes

jhonyy9 (jhonyy9):

and can you find now the 4th term too ?